A parallel plate capacitor has circular plates wit

Please show work
Answer #1

C = epsilion * A/d

A is area of plate

d is distance between plates

C = 8.85*10^-12 * pi*0.1*0.1/ 0.001

C = 2.78*10^-10 F

Q = CV

Q = 2.78*10^-10*12

Q = 3.33*10^-9 C

option A is correct